# Charged excitations made neutral: N-centered ensemble density functional theory of Fukui functions

### v1

@@ Abstract
We introduce an in-principle exact working equation to compute Fukui functions within $N$-centered ensemble DFT. It avoids the kernel derivative discontinuity problem of DFT for fractional number of electrons, whose contribution is recovered through weight-derivatives of the ensemble density functional potential. We explore practical strategies to compute its contribution by recycling ground-state density-functional approximations dressed with a weight-dependent scaling function. We also show interpolating between known limits of the ensemble functional and learning from uniform density profiles is a very effective strategy.
@@
As chemistry and condensed-matter physics is first and foremost dictated by the response of electrons to perturbation (from other electrons and nuclei, light, applied electric field and many more), they are the focus of properties descriptors. The advent of Density Functional Theory (DFT), an in-principle exact framework in which every property of the system follows from the electronic density, brought a solid foundation from which sharp definitions of insightful concepts – such as electronegativityParr (1980) and hardnessParr and Pearson (1983) – emerged. Further analysis of increasingly local responses of the electronic energy to changes in the external potential and number of electrons bloomed into the field of Conceptual Density Functional TheoryGeerlings et al. (2003); Geerlings and De Proft (2008) (CDFT). In turn, their computation from corresponding functional derivatives of the (to-be-approximated) universal functional proved instrumental in understanding what information practical Density Functional Approximations (DFA) should incorporate: By focusing efforts on reproducing more relevant properties, DFA design can (hopefully) find shortcuts on the meandering path towards the exact functionalGould (2017). This is a contribution of the present work.
@@
We will focus more specifically on the Fukui function, introduced by Parr and YangParr and Yang (1984) as a tool to understand the relative reactivity of different sites in a moleculeAyers and Levy (2000). It is defined as the derivative of the electronic density with respect to the electron number, a somewhat puzzling concept for isolated moleculesNguyen-Dang et al. (1982) which must have an integer number of electrons, but very natural for atoms or fragments within a bigger molecular arrangement resulting in partial charges on each subunitBader (1990). Based on the reasoning that an approaching reagent will prefer the site most favorable to addition/removal of an electron for a nucleophilic/electrophilic attack, the corresponding Fukui function will reach a higher value the more reactive a site isAyers and Levy (2000). It offers a (sometimes criticalBartolotti and Ayers (2005)) improvement over frontier molecular orbital theoryFukui et al. (1952, 1954); Fukui (1982) by inclusion of orbital relaxation effects.
@@
Interpretation and computation of derivatives with respect to the number of electrons analytically, i.e. connecting them to the $N$-electron Kohn Sham system of DFT and the Hartree-exchange-correlation (Hxc) density functional, is much more insightful (and more accurate) than their finite-differences evaluation from separate calculations on the $N$ and NN-1 electron systemsYang et al. (2012). To that end, extension of DFT to fractional electron numbers by Perdew, Parr, Levy and Balduz (PPLB) showed the generalized energy functional possesses derivative discontinuities when crossing an integer number of electronsPerdew et al. (1982), a behavior which in turn affects properties obtained through differentiating itPerdew et al. (1982); Mori-Sánchez et al. (2008); Cohen et al. (2008a, b); Stein et al. (2010); Zheng et al. (2011); Cohen et al. (2011); Perdew et al. (2017). Accurate predictions depend on inclusion of discontinuous shifts of the Hxc fields at integer electron number thresholds, which practical (local and semi-local) DFAs lack. The most well-known case is the missed contribution of Hxc potential derivative discontinuities resulting in systematic underestimation of fundamental gaps in solidsPerdew et al. (2017), also relevant for molecular dissociation and charge transferHodgson et al. (2017); Kraisler et al. (2021). It is less widely appreciated that similar discontinuities also appear in the Hxc kernel, responsible for aforementioned orbital relaxation effects, making computation of the Fukui function within PPLB challengingHellgren and Gross (2012). Constraints on KS orbital densities arising from piecewise-linearity requirement of the physical density were recently investigatedKraisler (2025), aiming to identify and correct density-driven errors in DFT calculations. The HOMO dependence on its fractional occupation proved to be intricate, invalidating the frozen approximation and even linear behavior in simple atomic systems. In some cases, it was found to not even be two-point-Taylor-expandable, with preliminary analysis suggesting the divergence of Taylor coefficients stems from approximate functionals being used.
@@
The recently introduced $N$-centered extension of ensemble density functional theorySenjean and Fromager (2018); Hodgson et al. (2021); Cernatic et al. (2022) (Nc EDFT) is an alternative exact approach to charged electronic excitations that shed a new light on the discontinuity problem. While PPLB considers an ensemble of $N$ and N1N-1 electons pure states with relative weights constraining the obtained ensemble density, the Nc ensemble density always integrates to the central number of electrons $N$ by construction, no matter the weights (vide infra). The ensemble density and weights become independent variables of the theory, which transform the elusive energy derivative discontinuity contributions into weight-derivatives of the Hxc ensemble functional. This offers a promising way to ”recycle” local KS DFA lacking derivative discontinuities. It also provides a considerable gain in scope as it can describe any charged or neutral excitation within a unified pictureCernatic et al. (2024), and can be adapted to treat open systemsSenjean and Fromager (2020). We can benefit from progress on DFA development for EDFT in recent yearsGould and Pittalis (2019); Fromager (2020); Gould et al. (2020, 2023); Gould and Pittalis (2024); Gould et al. (2025a, b). Moreover, the introduction by the Authors of in principle exact expressions for neutrally-excited state densitiesFromager (2025) and responsesDupuy and Fromager (2025) within an ensemble makes computation of properties beyond excitation energies a timely pursuit.
@@
In this Letter, we derive an exact $N$-centered EDFT of Fukui functions and discuss the fundamental and practical implications of our findings in the description of second (and possibly higher)-order density-functional Hxc derivatives. We then investigate two strategies to build a practical EDFAs and study their performance on the Hubbard dimer model systemCarrascal et al. (2015) as proof of principle. First, we will show how dressing ground-state KS DFAs with a weight-dependent scaling function allows to capture the weight-derivative term. Then, we will demonstrate the performance of an EDFA built by interpolating between known low and high correlation limits of the Hxc kernel, with additional constraint from the ground-state uniform electron gas limit.
@@
$n_{\Psi_{\nu}}=\frac{N_{\nu}}{N}n^{\bm{\xi}}+\sum_{\lambda>0}(\delta_{\lambda\nu}-\frac{N_{\nu}}{N}\xi_{\lambda})\frac{\partial n^{\bm{\xi}}}{\partial\xi_{\lambda}}$.
According to the KS mapping of Eq. (3), it leads to the more explicit expression (where we dropped the $\mathbf{r}$ dependence, for the sake of compactness),
@@
where we use the shorthand notation $\chi\star f:\mathbf{r}\mapsto\int d\mathbf{r}^{\prime}\chi(\mathbf{r},\mathbf{r}^{\prime})f(\mathbf{r}^{\prime})$ and $\chi^{\bm{\xi}}_{s}$ is the non-interacting KS analog of the Nc ensemble density-density linear response function $\chi^{\bm{\xi}}:\,(\mathbf{r},\mathbf{r}^{\prime})\mapsto\chi^{\bm{\xi}}(\mathbf{r},\mathbf{r}^{\prime})=\delta n^{\bm{\xi}}(\mathbf{r})/\delta v(\mathbf{r}^{\prime})$. Note that the two response functions are related through the following Dyson equation,
@@
Note that Eq. (8) generalizes Eq. (104) of Ref. 41 to charged excited states. As it applies universally to any $N_{\nu}$-electron ground or excited state ($N_{\nu}=N,N\pm 1,N\pm 2,\ldots$), it can be (straightforwardly) applied to the calculation of physical (interacting) Fukui functions $f_{\kappa}={\theta}_{\kappa}(n_{\Psi_{\kappa}}-n_{\Psi_{0}})$, for which $N_{\kappa}\neq N$ and ${\theta}_{\kappa}={\rm sgn}(N_{\kappa}-N)$, thus leading to the following exact relation between $f_{\kappa}$ and its non-interacting KS analog $f^{\bm{\xi}}_{s,\kappa}={\theta}_{\kappa}(n_{\Phi^{\bm{\xi}}_{\kappa}}-n_{\Phi^{\bm{\xi}}_{0}})$,
@@
Interestingly, in regular (PPLB) DFT, this relation reduces to the first term on the right-hand side (the second and third terms simply do not exist in PPLB), which is then dependent on the aforementioned discontinuous Hxc kernel for accurate predictions Hellgren and Gross (2012). It appears that, within the Nc formalism, the impact of discontinuities on the calculation of Fukui functions is described explicitly through ensemble weight derivatives, as readily seen from the last term on the right-hand side of Eq. (9). Thus, we generalize a feature that was already known for the evaluation of ionization potentials and electron affinities Cernatic et al. (2024). This equivalence can be made even more transparent by using the invariance of Eq. (9) under symmetric shifts
@@
Eq. (12), which is our second and central formal result, establishes the equivalence between the discontinuities in the Hxc kernel and the weight derivatives of the ensemble Hxc density-functional potential (see the first term on the right-hand side). The clarity of the Nc EDFT picture is made manifest in the fact that no assumption about the asymptotic behavior of the density needs to be invoked Hellgren and Gross (2012) to rationalize and evaluate the shift. This is relevant for analysis of periodic systems, where Fukui functions can rationalize surface reactivity Barrera et al. (2025).
Practical strategies for EDFA design – By exploring the performance of EDFAs for computing Fukui functions through Eq. (9), we will discuss what a good approximation to the ensemble Hxc functional is, gather insights on the interplay of density-dependence and weight-dependence, and particularly, how Nc EDFT could possibly recycle semi-local standard DFAs whose lack of discontinuous behavior impeded PPLB. We do so by applying the Nc formalism to the Hubbard Dimer as it allows to study all regimes of electronic correlation (controlled by the parameter $U/t$) while the Hxc functional and its derivatives can be computed exactlyCarrascal et al. (2015). Note the strategies are general, and applicable to Ab Initio systems. We will focus on charged excitations between different electronic ground states, considering an ensemble of the $N=2$-electron state and the ”neighboring” $N\pm 1$-electron states with weights $\bm{\xi}=(\xi_{-},\xi_{+})$. In all results shown, we evaluate density functionals at the exact ensemble density. The ensemble response function $\chi^{\bm{\xi}}$ will be computed from Eq. (5) using the exact KS response.
The opportunity to employ regular (ground-state) DFAs arises from noticing that Eq.(9) holds even in the zero weights limit, where the ensemble Hxc kernel, response and density become those of regular $N$-electron ground-state DFT:
@@
$f_{s,+}(\mathbf{r})$ is the lowest unoccupied KS molecular orbital (LUMO) in the $N$-electron ground state and $f_{s,-}(\mathbf{r})$ the highest occupied one (HOMO). Note $n_{0}=n^{\bm{\xi}=0}=n_{\Psi_{0}}$.

### v2

@@ Abstract
An in-principle exact working equation to compute electronic affinity and ionization Fukui functions is derived within the $N$-centered (Nc) ensemble extension of density functional theory (DFT). It circumvents the kernel derivative discontinuity problem of DFT for fractional electron numbers, whose contribution is recovered through weight derivatives of the ensemble density functional potential. Thus, it allows for the design of alternative and effective approximations, such as the weight-dependent scaling of regular functionals or the interpolation between known limits of Nc ensembles.
@@
As chemistry and condensed-matter physics is first and foremost dictated by the response of electrons to perturbation (from other electrons and nuclei, light, applied electric field and many more), they are the focus of properties descriptors. The advent of Density Functional Theory (DFT), an in-principle exact framework in which every property of the system follows from the electronic density, brought a solid foundation from which sharp definitions of insightful concepts – such as electronegativity Parr (1980) and hardness Parr and Pearson (1983) – emerged. Further analysis of increasingly local responses of the electronic energy to changes in the external potential and number of electrons bloomed into the field of Conceptual Density Functional Theory Geerlings et al. (2003); Geerlings and Proft (2008) (CDFT). In turn, their computation from corresponding functional derivatives of the (to-be-approximated) universal functional proved instrumental in understanding what information practical Density Functional Approximations (DFA) should incorporate: By focusing efforts on reproducing more relevant properties, DFA design can (hopefully) find shortcuts on the meandering path towards the exact functional Gould (2017). This is a contribution of the present work.
@@
We will focus more specifically on the Fukui function, introduced by Parr and Yang Parr and Yang (1984) as a tool to understand the relative reactivity of different sites in a molecule Ayers and Levy (2000). It is defined as the derivative of the electronic density with respect to the electron number, a somewhat puzzling concept for isolated molecules Nguyen-Dang et al. (1982) which must have an integer number of electrons, but very natural for atoms or fragments within a bigger molecular arrangement resulting in partial charges on each subunit Bader (1990). Based on the reasoning that an approaching reagent will prefer the site most favorable to addition/removal of an electron for a nucleophilic/electrophilic attack, the corresponding Fukui function will reach a higher value the more reactive a site is Ayers and Levy (2000). It offers a (sometimes critical Bartolotti and Ayers (2005)) improvement over frontier molecular orbital theory Fukui et al. (1952, 2006); Fukui (1982) by inclusion of orbital relaxation effects.
@@
Interpretation and computation of derivatives with respect to the number of electrons analytically, i.e. connecting them to the $N$-electron Kohn Sham (KS) system of DFT and the Hartree-exchange-correlation (Hxc) density functional, is much more insightful (and more accurate) than their finite-differences evaluation from separate calculations on the $N$ and N1N-1 electron systems Yang et al. (2012). To that end, extension of DFT to fractional electron numbers by Perdew, Parr, Levy and Balduz (PPLB) showed the generalized energy functional possesses derivative discontinuities when crossing an integer number of electrons Perdew et al. (1982), a behavior which in turn affects properties obtained through differentiating it Perdew et al. (1982); Mori-Sánchez et al. (2008); Cohen et al. (2008a, b); Stein et al. (2010); Zheng et al. (2011); Cohen et al. (2011); Perdew et al. (2017). Accurate predictions depend on inclusion of discontinuous shifts of the Hxc fields at integer electron number thresholds, which practical (local and semi-local) DFAs lack. The most well-known case is the missed contribution of Hxc potential derivative discontinuities resulting in systematic underestimation of fundamental gaps in solids Perdew et al. (2017), also relevant for molecular dissociation and charge transfer Hodgson et al. (2017); Kraisler et al. (2021). It is less widely appreciated that similar discontinuities also appear in the Hxc kernel, responsible for aforementioned orbital relaxation effects, making computation of the Fukui function within PPLB challenging Hellgren and Gross (2012). Constraints on KS orbital densities arising from piecewise-linearity requirement of the physical density were recently investigated Kraisler (2025), aiming to identify and correct density-driven errors in DFT calculations. The highest occupied molecular orbital (HOMO) dependence on its fractional occupation proved to be intricate, invalidating the frozen approximation and even linear behavior in simple atomic systems. In some cases, it was found to not even be two-point-Taylor-expandable, with preliminary analysis suggesting the divergence of Taylor coefficients stems from approximate functionals being used.
@@
The recently introduced $N$-centered extension of ensemble density functional theory Senjean and Fromager (2018); Hodgson et al. (2021); Cernatic et al. (2022) (Nc EDFT) is an alternative exact approach to charged electronic excitations that shed a new light on the discontinuity problem. While PPLB considers an ensemble of $N$ and N1N-1 electrons pure states with relative weights constraining the obtained ensemble density, the Nc ensemble density always integrates to the central number of electrons $N$ by construction, no matter the weights (vide infra). The ensemble density and weights become independent variables of the theory, which transform the elusive energy derivative discontinuity contributions into weight-derivatives of the Hxc ensemble functional. This offers a promising way to ”recycle” (semi-) local DFAs lacking derivative discontinuities Loos and Fromager (2020). It also provides a considerable gain in scope as it can describe any charged or neutral excitation within a unified picture Cernatic et al. (2024), and can be adapted to treat open systems Senjean and Fromager (2020). The theory can also benefit from progress made in recent years in ensemble DFA development Gould and Pittalis (2019); Fromager (2020); Gould et al. (2020, 2023); Gould and Pittalis (2024); Gould et al. (2026, 2025). Moreover, the very recent derivation by the authors of exact expressions for neutrally-excited state densities Fromager (2025) and their responses Dupuy and Fromager (2025) within an ensemble makes computation of properties beyond excitation energies a timely pursuit.
@@
In this Letter, we derive an exact $N$-centered EDFT of Fukui functions and discuss the fundamental and practical implications of our findings in the description of second (and possibly higher)-order density-functional Hxc derivatives. We then investigate two strategies to build a practical EDFA and study their performance on the Hubbard dimer model system Carrascal et al. (2015) as proof of principle. First, we will show how dressing the regular ground-state Hxc functional with a weight-dependent scaling function allows to capture the weight-derivative term. Then, we will demonstrate the performance of an EDFA built by interpolating between known low and high correlation limits of the Hxc ensemble density functional.
@@
where we use the shorthand notation $\chi\star f:\mathbf{r}\mapsto\int d\mathbf{r}^{\prime}\chi(\mathbf{r},\mathbf{r}^{\prime})f(\mathbf{r}^{\prime})$, and $\chi^{\bm{\xi}}_{s}$ is the non-interacting KS analog of the Nc ensemble density-density linear response function $\chi^{\bm{\xi}}:\,(\mathbf{r},\mathbf{r}^{\prime})\mapsto\chi^{\bm{\xi}}(\mathbf{r},\mathbf{r}^{\prime})=\delta n^{\bm{\xi}}(\mathbf{r})/\delta v(\mathbf{r}^{\prime})$. Note that the two response functions are related through the following Dyson equation,
@@
Note that Eq. (8) generalizes Eq. (104) of Ref. 42 to charged excited states. As it applies universally to any $N_{\nu}$-electron ground or excited state ($N_{\nu}=N,N\pm 1,N\pm 2,\ldots$), it can be (straightforwardly) applied to the calculation of physical (interacting) Fukui functions $f_{\kappa}={\theta}_{\kappa}(n_{\Psi_{\kappa}}-n_{\Psi_{0}})$, for which $N_{\kappa}\neq N$ and ${\theta}_{\kappa}={\rm sgn}(N_{\kappa}-N)$, thus leading to the following exact relation between $f_{\kappa}$ and its non-interacting KS analog $f^{\bm{\xi}}_{s,\kappa}={\theta}_{\kappa}(n_{\Phi^{\bm{\xi}}_{\kappa}}-n_{\Phi^{\bm{\xi}}_{0}})$,
@@
Interestingly, in regular (PPLB) DFT, this relation reduces to the first term on the right-hand side (the second and third terms simply do not exist in PPLB), which is then dependent on the aforementioned discontinuous Hxc kernel for accurate predictions Hellgren and Gross (2012). It appears that, within the Nc EDFT formalism, the impact of discontinuities on the calculation of Fukui functions is described explicitly through ensemble weight derivatives, as readily seen from the last term on the right-hand side of Eq. (9). Thus, we generalize a feature that was already known for the evaluation of ionization potentials and electron affinities Cernatic et al. (2024). This equivalence can be made even more transparent by using the invariance of Eq. (9) under symmetric shifts in the ensemble Hxc kernel,
@@
Eq. (12), which is our second and central formal result, establishes the equivalence between the discontinuities in the Hxc kernel and the weight derivatives of the ensemble Hxc density-functional potential (see the first term on the right-hand side). The clarity of the Nc EDFT picture is made manifest in the fact that no assumption about the asymptotic behavior of the density needs to be invoked Hellgren and Gross (2012) to rationalize and evaluate the shift. This is relevant for analysis of periodic systems, where Fukui functions can rationalize surface reactivity Barrera et al. (2025), for example.
Practical strategies for Nc EDFA design – The opportunity to employ regular (ground-state) DFAs in this context arises from noticing that Eq. (9) holds even in the zero-weight limit of the theory, where the ensemble Hxc kernel, response function and density become those of regular $N$-electron ground-state DFT. If we focus more specifically on the ground-state theory of a single electron addition ($\kappa\equiv+$) or removal ($\kappa\equiv$-), it reads